OpenStudy (anonymous):

How would I put cos(u^2) in to terms of x using u sub. Could I....

OpenStudy (anonymous):

say u=x^2 and du= 2x dx and then say that the integral cos(u^2)du is equal to cos(x)*2x dx

OpenStudy (anonymous):

switch that u^2=x

OpenStudy (anonymous):

I'm not sure I follow what you're trying to do? Typically you'll have something like \[\int u\ cos(u^2)\ du\]

OpenStudy (anonymous):

the question is integrate cos(u^2)du from cos(x) to 5x

OpenStudy (anonymous):

My mistake, Find the derivative of the following y= the integral from cos(x) to 5x of cos(U^2)du

OpenStudy (anonymous):

Oh, whew.. Much better ;p

OpenStudy (anonymous):

So I assume you put u into terms of x, then integrate, then evaluate, then take the derivative?

OpenStudy (anonymous):

Are you taking the derivative with respect to x or with respect to u?

OpenStudy (anonymous):

x I think.

OpenStudy (anonymous):

It doesn't say?

OpenStudy (anonymous):

It should say.

OpenStudy (anonymous):

I will put in the equation as stated

OpenStudy (anonymous):

Find the derivative of the following function: \[y=\int\limits_{cosx}^{5x} \cos(u^2)du\]

OpenStudy (anonymous):

so I decided to let x=u^2 and du=2x so the equation becomes integral of cos(x)*2xdx

OpenStudy (anonymous):

I don't think you need to do that. I feel as though they want you to rely on the fundamental theorem of calculus here..

OpenStudy (anonymous):

Wait so I can just take the integral of cos(u^2) and then eval for 5x and cos x?

OpenStudy (anonymous):

and then take the derivative of the evaluated function?

OpenStudy (anonymous):

Oh, wait, sure you can.. just requires a u-sub, then parts. Sorry.

OpenStudy (anonymous):

\[\int\limits 2xcos(x) dx \rightarrow 2xsin(x)+2\cos(x)\]

OpenStudy (anonymous):

I dunno where to go with it. The answer is sin(x)*cos(cos^2x)+5cos(25x^2)

OpenStudy (anonymous):

Ah forgot to change the limits.

OpenStudy (anonymous):

You have to do parts again on 2xsin(x) so it repeats and gets you back to 2xcos(x)

OpenStudy (anonymous):

It doesn't repeeat does it U=2x du =2 v=sin(x) dv = cos(x)

OpenStudy (anonymous):

Oh, you're right.

OpenStudy (anonymous):

Wait does the question actually ask? \[d/du \int\limits_{\cos(x)}^{5x} \cos(u^2)du\]

OpenStudy (anonymous):

Do the integral and the derivative cancel?

OpenStudy (anonymous):

Wait, what did you use for the original u-sub?

OpenStudy (anonymous):

u^2=x

OpenStudy (anonymous):

\[u^2 = x \implies 2u\ du = dx \implies du = \frac{1}{2\sqrt{x}}dx\]

OpenStudy (anonymous):

So I don't think you're using the right integral. \[\int cos(u^2)du = \int {cos(x) \over 2\sqrt{x}}dx\]

OpenStudy (anonymous):

Probably I was using du=2xdx

OpenStudy (anonymous):

I can see why it would be 1/2x I don't see how it is root x

OpenStudy (anonymous):

Because \(u^2 = x \implies u = \sqrt{x}\) So \[2u\ du = dx \implies du = \frac{1}{2u}dx = \frac{1}{2\sqrt{x}}dx\]

OpenStudy (anonymous):

I think we need to use the fundamental theorem. And I think it'll make this much easier..

OpenStudy (anonymous):

Oh I see.

OpenStudy (anonymous):

Lets see here. Lets assume that \[F(u) = \int cos(u^2) du\] \[\implies \int\limits_{\cos(x)}^{5x} \cos(u^2)du = F(5x) - F(cos(x))\] Right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Ok, well if \[F(u) = \int cos(u^2) du\] Then\[\frac{d}{du}F(u) = cos(u^2) \]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

I see so the integral and the derivative cancel?

OpenStudy (anonymous):

differential*

OpenStudy (anonymous):

Yes. That's the fundamental theorem of calculus. Integration and derivatives are inverses of eachother (plus or minus a constant).

OpenStudy (anonymous):

So the question is just asking cos(5x^2)-cos(cos(x^2)

OpenStudy (anonymous):

So that means that \[\frac{d}{dx} \int\limits_{\cos(x)}^{5x} \cos(u^2)du =\frac{d}{dx}[ F(5x) - F(cos(x))]\]\[= cos((5x)^2)( \frac{d}{dx}[5x]) - cos(cos^2(x))(\frac{d}{dx}[cos(x)])\]\[=?\]

OpenStudy (anonymous):

Oh snap. 5cos(25x^2)+sin*cos(cos^2(x)) which is the answer.

OpenStudy (anonymous):

So just have to remember that the derivative undoes the integral, then be sure to use the chain rule since your limits are functions of another variable.

OpenStudy (anonymous):

That's it great thanks.